In a multiple regression analysis, two independent variables are considered, and the sample size is 26. The regression coefficients and the standard errors are as follows.

b1 = 1.468 Sb1 = 0.89
b2 = ?1.084 Sb2 = 0.88

Conduct a test of hypothesis to determine whether either independent variable has a coefficient equal to zero. Would you consider deleting either variable from the regression equation? Use the 0.05 significance level. (Negative amounts should be indicated by a minus sign. Round your answers to 3 decimal places.)

H0: ?1 = 0 H0: ?2 = 0
H1: ?1 ? 0 H1: ?2 ? 0
H0 is rejected if t < ?2.069 or t > 2.069
Answer #1

For X1:

Degree of freedom: df=n-2=24

The test statistics is

t=\frac{b_{1}-\beta_{1}}{sb_{1}}=\frac{1.468}{0.89}=1.649

Since test statistics is not lie in the rejection region so we fail to reject the null hypothesis.

-----------------

For X2:

Degree of freedom: df=n-2=24

The test statistics is

t=\frac{b_{2}-\beta_{2}}{sb_{2}}=\frac{-1.084}{0.88}=-1.232

Since test statistics is not lie in the rejection region so we fail to reject the null hypothesis.

That is both variables are not significant to model.